OpenStudy (anonymous):

(x+4)(x+2) multiply and simplify

OpenStudy (anonymous):

(x^2+6x+8)

OpenStudy (anonymous):

x^2+6x+8

OpenStudy (anonymous):

(x+4)(x+2)=x^2+4x+2x+8=x^2+6x+8